I got 142 marks in IIT JEE and my Intemediate score is 979 out of 1000 for Andhra pradesh board. What is my expected AIR ?
 

Grade:12

2 Answers

Saurabh Kumar
8 years ago
Pls Use askiitians.com rank predictor....
As per 60 – 40 ratio, you would have a decent AIR, ..
thabitha
8 years ago
Thanks for the answer.  But i am confused in using the askiitians.com rank predictor.
For AP board, Inter 1st year and 2nd year complete score is required. Rank predictor is expecting marks in Physics, chemistry and Maths -
Physics : Total marks = 60 (Inter 1st year ) + 60 (Inter IInd year) = 120, my score is 118.
Chemistry : Total  marks= 60 (Inter 1st year ) + 60( Inter IInd year) = 120, my score is 116.
Maths  : Total  marks= 150 (Inter 1st year ) + 150( Inter IInd year) = 300, my score is 300.
Rank predictor is not accepting to enter more than 100 for Maths.  
 
Please help. Will my AIR is below 10K ?